OpenStudy (abbycross167):

Can someone please show me how to answer this question? I will fan and give a medal.... Question~ Solve the systems using the elimination method. y = 2x - 1 y = -x + 7

OpenStudy (anonymous):

Multiply each side of the second equation by -1. Add the two equations together. The y terms will subtract out.

OpenStudy (abbycross167):

so the second equation would look like this after I multiply -- y=1+-7?

OpenStudy (abbycross167):

@robtobey

OpenStudy (anonymous):

What do u think

OpenStudy (abbycross167):

umm what? I think the equation is supposed to look like y=1+-7 but I'm not sure

OpenStudy (mathmale):

Please actually follow Tobey's instructions: "Multiply each side of the second equation by -1. Add the two equations together. The y terms will subtract out." What's left? You will hav e enuf info there to be able to solve for x.

OpenStudy (mathmale):

abby: "I think the equation is supposed to look like y=1+-7" is either not a function or not a very good one. Where's your independent variable?

OpenStudy (anonymous):

okay first u rearrange them it should look like this -2x+y=-1 x+y=7

OpenStudy (anonymous):

-2x+y=-1-----------------X1 (Multiply by 1) x+y=7--------------------X2 (Multiply by 2)

OpenStudy (anonymous):

i think u can take it from here

OpenStudy (mathmale):

Abby: Take that as a good challenge. Show your stuff!